A reversible compression of 1 mol of an ideal gas
Answer #1

As,

PV=NRT

as no. of moles are constant

So

PV/T=Constant

(T/P) *constant=V

also

PV1.55 = constant

P((T/P) *constant)^1.55=constant

P^(-0.55)*(T)^(1.55)= constant

P2/P1 = T2/T1

P2 = T2/T1 * P1

AS, P1 = 1

SO,

P2 = 950/400

P2 = 2.375 Bars

Cp- Cv= R

CP/R = 3.85 + 0.57 x10-3 T

From first law of Thermodyanamics

the heat transferred is calculated as.

Q = W + dU

BY putting values In above equation,

We get the heat required for the process